OpenStudy (anonymous):

write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in the standard form, ax^2 + bx + c = 0. solution: 9, only

OpenStudy (anonymous):

Its bascially a formula

OpenStudy (anonymous):

since only one root i.e. 9 is given, it means that the quadratic polynomial has two equal roots ie.e 9 and 9 This means that equation will have factors (x-9)(x-9) so the equation is x^2 -9x -9x +81 so the required equation is x^2 - 18x + 81

OpenStudy (anonymous):

Sorry, required equation is x^2 - 18x + 91 = 0

OpenStudy (anonymous):

but harkirat its a case

OpenStudy (anonymous):

91?

OpenStudy (anonymous):

what do u mean narender, pls explain...

OpenStudy (anonymous):

3u/u^2-1 + u/u-1, simplify…

OpenStudy (anonymous):

sorry again!!! required equation is x^2 - 18x + 81 = 0 PS: I am attending to too many questions simultaneously, so ....

OpenStudy (anonymous):

thank you Harkirat

OpenStudy (anonymous):

can you please solve the above problem too Hari?

OpenStudy (anonymous):

3u/u^2-1 + u/u-1, simplify…

OpenStudy (anonymous):

is it.. 3u u -------- + ------ ???? u^2 - 1 u + 1

OpenStudy (anonymous):

u/ u-1 yes

OpenStudy (anonymous):

ok, this is how we do it..... We use the identity \[a ^{2} - b ^{2} = (a+b)(a-b)\] 3u u 3u u 3u u -------- + ------ = ----------- + ------ = ----------- + ------- u^2 - 1 u - 1 u^2 - 1^2 u-1 (u+1)(u-1) u-1 3u + u (u+1) 3u + u^2 + u u^2 + 4u = ---------------- = -------------- = ------------ (u+1) (u-1) (u+1)(u-1) (u+1)(u-1)

OpenStudy (anonymous):

I don't know if it can be simplified any further..........

OpenStudy (anonymous):

perfect it can be simplified to u(u+4)/ (u-1)(u+1)

OpenStudy (anonymous):

correct, actually i am dead tired now and am logging off..... It is right now 2 AM in India !!!!!!

OpenStudy (anonymous):

where in INdia are you from

OpenStudy (anonymous):

Its capital- Delhi !!!!

OpenStudy (anonymous):

n where r u located??

OpenStudy (anonymous):

I am originally from Hyderabad

OpenStudy (anonymous):

I live in St Louis USA

OpenStudy (anonymous):

i see..............

OpenStudy (anonymous):

home schooling???

OpenStudy (anonymous):

what do you do for a living Hari?

OpenStudy (anonymous):

Kinda

OpenStudy (anonymous):

i work in a private company and teach some students after coming back from office....

OpenStudy (anonymous):

the thing on your ID, is that the golden temple?

OpenStudy (anonymous):

You are so good both as a math tutor and a human being

OpenStudy (anonymous):

absolutely right, I had visited it in January and this is a photo I took then...

OpenStudy (anonymous):

are you into photography?

OpenStudy (anonymous):

thx for saying that.....

OpenStudy (anonymous):

Yeah, i have been using cameras for some 15-17 years now but mostly family photos and all wish had more liesurely life to shoot nature etc...

OpenStudy (anonymous):

do you use telephoto lenses?

OpenStudy (anonymous):

I have automatic cameras with zoom

OpenStudy (anonymous):

I have a canon 7D and 70-200 telephoto, and a 100 macro, and 10 fish-eye

OpenStudy (anonymous):

I love photography!

OpenStudy (anonymous):

I guess u r a keen photographer...

OpenStudy (anonymous):

ok, it is a good hobby and can become a profession too...

OpenStudy (anonymous):

yes I am :)

OpenStudy (anonymous):

are you married?

OpenStudy (anonymous):

Yes, i hv two lovely daughters.....

OpenStudy (anonymous):

how old are they?

OpenStudy (anonymous):

I can send them some candy if you think its okay?

OpenStudy (anonymous):

one is 15 years old and the other is 9 years...

OpenStudy (anonymous):

they must be keeping you busy.

OpenStudy (anonymous):

Born and raised in Delhi?

OpenStudy (anonymous):

yes they do By the way what is yr real name and what standard r u in??

OpenStudy (anonymous):

yes we all have been born n brought up in Delhi..

OpenStudy (anonymous):

its getting late Hari, go get some sleep. Thank you once again for your time and help